OpenStudy (anonymous):

How do you get R= S/0.75 from S= R-0.25R (inverse function)

OpenStudy (phoenixfire):

R-0.25R=0.75R S=0.75R --> R=S/0.75 dividing both sides by 0.75.

OpenStudy (anonymous):

no the original equation was S=R-0.25R. I want the inverse function and i wanna know how

OpenStudy (phoenixfire):

The inverse function is the function that reverses the mapping. so if you have a function S=R-0.25R that's the same as S=0.75R, which is a straight line. For every value of R it will produce a value of S. The inverse function is one where you have every value of S produce a value for R. So in this case, all you need to do, is rearrange the equation to get R= something to do with S.

OpenStudy (anonymous):

Okay I got it now. Thanks